OpenStudy (anonymous):

How do you simplify lny+c=x-2ln|2+x| to y=Ce^x [(2+x)^-2] ?

OpenStudy (anonymous):

http://www.calcchat.com/book/Calculus-8e/ Chapter 6 Section R Exercise 21

OpenStudy (turingtest):

subtract c from both sides and raise e to the power of both sides, then use a few log identities

OpenStudy (anonymous):

so then y= -C+(e^x) - (e^2ln(2+x))?

OpenStudy (turingtest):

yes, now use \[a\ln x=\ln x^a\] on the middle term

OpenStudy (turingtest):

should be y= -C+(e^x) - (e^(-2ln|2+x|)

OpenStudy (anonymous):

y=c+ e^x-(2+x)^2

OpenStudy (anonymous):

negative 2???

OpenStudy (turingtest):

should be y= -C+(e^x) + (e^(-2ln|2+x|)

OpenStudy (turingtest):

well it's easier if you pit the -2 in the exponent

OpenStudy (turingtest):

put*

OpenStudy (turingtest):

oh heck I'm busy and I gave you wrong info

OpenStudy (anonymous):

ohh you just put the negative on top rather than leaving it ~ oh ok so then y= c+e^x+(2+x)^-2

OpenStudy (anonymous):

so then how does that get to be multiplied with both C and e^x?

OpenStudy (turingtest):

lny+c=x-2ln|2+x| to y=Ce^x [(2+x)^-2] y=e^(x-2ln|2+x|-c) y=e^x * e^{-2ln|2+x|) * e^-c

OpenStudy (turingtest):

the terms need to be multiplied, not added :P

OpenStudy (turingtest):

sorry gotta get dinner, brb

OpenStudy (anonymous):

kay thnx

OpenStudy (turingtest):

ok, so do you follow up to lny+c=x-2ln|2+x| to y=Ce^x [(2+x)^-2] y=e^(x-2ln|2+x|-c) y=e^x * e^(-2ln|2+x|) * e^-c ?

OpenStudy (turingtest):

in LaTeX\[\ln y+c=x-2\ln|2+x|\\y=e^{x-2\ln|x+2|-c}=e^xe^{-2\ln|x+2|}e^{-c}\]